OpenStudy (anonymous):

solve using the principle of zero products (10t + 7)(3t-9) = 0

OpenStudy (anonymous):

(3t-9) = 0 t=9/3 t=3 (10t + 7)=0 t=-7/10

OpenStudy (anonymous):

so the answer is -7/10, and 3 correct